### v2

# RL of Thoughts: Navigating LLM Reasoning with Inference-time Reinforcement Learning

Abstract
Despite rapid advancements in large language models (LLMs), the token-level autoregressive nature constrains their complex reasoning capabilities. To enhance LLM reasoning, inference-time techniques, including Chain/Tree/Graph-of-Thought(s), successfully improve the performance, as they are fairly cost-effective by guiding reasoning through external logical structures without modifying LLMs’ parameters. However, these manually predefined, task-agnostic frameworks are applied uniformly across diverse tasks, lacking adaptability. To improve this, we propose RL-of-Thoughts (RLoT), where we train a lightweight navigator model with reinforcement learning (RL) to generate task-adaptive logical structures at inference time, enhancing LLM reasoning. Specifically, we design five basic logic blocks from the perspective of human cognition. During the reasoning process, the trained RL navigator dynamically selects the suitable logic blocks and combines them into task-specific logical structures according to problem characteristics. Experiments across multiple reasoning benchmarks (AIME, MATH, GPQA, etc.) with multiple LLMs (GPT, Llama, Qwen, and DeepSeek) illustrate that RLoT outperforms established inference-time techniques by up to 13.4%. Remarkably, with less than 3K parameters, our RL navigator is able to make sub-10B LLMs comparable to 100B-scale counterparts. Moreover, the RL navigator demonstrates strong transferability: a model trained on one specific LLM-task pair can effectively generalize to unseen LLMs and tasks. Our code is open-source at https://anonymous.4open.science/r/RL-LLM-Reasoning-1A30.
1 Introduction
Recent years have witnessed unprecedented advancements in large language models (LLMs), achieving remarkable success across diverse natural language tasks (Chang et al., 2024), including translation (Xu et al., 2024), semantic analysis (Lan et al., 2024b; a), and information retrieval (Hao et al., 2024). Despite these advancements, the inherent token-level autoregressive nature of LLMs poses a significant limitation for complex reasoning tasks (Zhao et al., 2023), such as solving mathematical problems (Ahn et al., 2024) or answering intricate questions (Zhuang et al., 2023). These tasks require sophisticated logical structures and long-term dependencies that go beyond the scope of simple sequential token prediction, leaving a considerable gap between current LLM capabilities and the demands of advanced reasoning applications.
Plentiful research has been devoted to enhancing LLM reasoning. On one hand, fine-tuning approaches attain substantial improvements on pretrained LLMs (Zhong et al., 2024; DeepSeek-AI et al., 2025; Team et al., 2025). However, these methods demand massive computational resources and large-scale datasets, being costly to implement. On the other hand, inference-time techniques, exemplified by Chain-of-Thought (Wei et al., 2022), Tree-of-Thoughts (Yao et al., 2023), and Graph-of-Thoughts (Besta et al., 2024), offer a lightweight alternative by enhancing reasoning through predefined external logical structures. While cost-effective, their logical structures rely on manual design and are task-agnostic, lacking the adaptability to diverse reasoning tasks.
Addressing such limitations in inference-time techniques presents significant challenges. First, reasoning tasks span various domains, including mathematics, STEM, commonsense, etc., where tasks in each domain exhibit diverse characteristics, making it infeasible to manually design logical structures specified for each task. Second, complex reasoning tasks often require multiple steps, where the problem-solving status evolves after each step, requiring dynamic adjustments to the logical structure for subsequent reasoning. Therefore, predefined logical structures fail to adapt to the changes, limiting their effectiveness in stepwise reasoning tasks. These challenges highlight the need for more adaptive inference-time techniques to handle reasoning tasks with diversity and dynamics.
[…]
In summary, the main contributions of this work include:
  • We propose RL-of-Thoughts (RLoT), an inference-time technique that leverages RL to adaptively construct task-specific logical structures, enhancing LLM reasoning.
  • We conduct extensive experiments to verify the effectiveness of our method to improve LLM reasoning across various tasks. Compatible with multiple widely known LLMs, RLoT outperforms established inference-time techniques by up to 13.4%.
[…]
2 Preliminaries
2.1 Large Language Models (LLMs)
Large language models (LLMs) are a class of advanced neural networks characterized by parameter scales up to billions, primarily trained through next-token prediction objectives. Given a sequence $\{w_{1},w_{2},...,w_{t-1}\}$, the models output $w_{t}$ to maximize the observation likelihood in the corpus as:
$\prod_{t=1}^{T}P(w_{t}|w_{1},w_{2},...,w_{t-1}).$ (1)
Recent advancements in LLMs, exemplified by architectures like the GPT series (Brown et al., 2020; Kalyan, 2023; Achiam et al., 2023), the Llama family (Touvron et al., 2023; Dubey et al., 2024), etc, have demonstrated remarkable proficiency across diverse natural language understanding and generation tasks, including semantic parsing, cross-lingual translation (Zhao et al., 2023; Chang et al., 2024). Meanwhile, extensive researches integrate inference-time techniques like Chain-of-Thought (CoT) (Wei et al., 2022) and Tree-of-Thoughts (ToT) (Yao et al., 2023) to enhance the multi-step reasoning capability of LLMs. On the other hand, fine-tuning strategies leverage Outcome Reward Models (ORM) and Process Reward Models (PRM) to optimize the reasoning process through reward-guided learning (Lightman et al., 2023; Wang et al., 2024c; Luo et al., 2024). These approaches address both structural limitations of auto-regressive decoding and the challenge of maintaining logical coherence in complex tasks (Xu et al., 2025).
2.2 Markov Decision Process (MDP)
A Markov Decision Process (MDP) provides the core framework for sequential decision-making problems. An MDP is mathematically defined by the tuple $(\mathcal{S},\rho,\mathcal{A},P,R)$, where $\mathcal{S}$ is the state space, and $\rho\in\Delta(\mathcal{S})$ represents the probability distribution over initial states, with $\Delta(\mathcal{S})$ being the set of all probability distributions over $\mathcal{S}$. The action space is denoted by $\mathcal{A}$. Given a specific action taken in a particular state, the state transition probability function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ and the reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ define the likelihood of transitioning between states and the reward associated with each action. At each time step $t$, the agent chooses an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$, receives a reward $r_{t}$, and transitions to the next state $s_{t+1}$. The agent’s objective in an MDP is to maximize the total accumulated reward over time, which is the sum of the discounted rewards obtained at each step. The cumulative reward at time step $t$ is expressed as $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}$, where $\gamma$ is the discount factor that weighs the significance of future rewards.
3 Methods
3.1 Overview
Figure 1: Framework of RL-of-Thoughts (RLoT). We train an RL agent as the navigator, which dynamically selects and combines basic logic blocks along the reasoning process, constructing task-specific logical structures for each task and thereby enhancing the LLMs’ ability to handle complex reasoning tasks.
In this paper, we propose to enhance LLM reasoning at inference time with RL, namely RL-of-Thoughts (RLoT). The overall framework of RLoT is illustrated in Figure 1. Specifically, we first model long-sequence reasoning as a Markov Decision Process (MDP), where we design the action space, state space, and state transition mechanism, making the sequence of decisions within an episode correspond to the generation of a logical structure for reasoning (Section 3.2). Within this MDP framework, we train an RL agent, referred to as the navigator model, which conducts sequential decision-making to construct task-specific logical structures, thereby enhancing LLMs’ capability to address complex reasoning tasks (Section 3.3).
3.2 LLM Reasoning as MDP
To leverage the sequential decision-making capability of RL for adaptively designing logical structures at inference time based on problem characteristics, we first model long-sequence reasoning as a specially designed MDP. Within this MDP, the sequence of decisions in an episode corresponds to the generation of a logical structure for reasoning. The designs of the state, action, reward, and state transition are as follows.
State. The state space is designed to capture how the current solving status of the task is after steps of reasoning, thereby supporting dynamic adjustments to the logical structure for subsequent reasoning. We employ a self-evaluation mechanism to extract concise and informative states during the reasoning process. Specifically, we prompt the LLM itself to evaluate the current reasoning steps from three major and seven detailed aspects, which are listed in Table 1.
Table 1: Aspects for state self-evaluation.
Major A: Correctness B: Complexity C: Completeness
Detailed A1: Correctness of modeling B1: Complexity to the final answer C1: Closeness to the final solution
A2: Clarity for further reasoning
B2: Alternative methods in
further reasoning
C2: Completeness within the step
A3: Correctness of calculation
For each detailed aspect, we require the LLM to assign a score from 1 to 3, which is then aggregated to form the state of the MDP. This approach summarizes complex reasoning steps into low-dimensional states, offering a comprehensive overview of the changing problem-solving status during the reasoning process, facilitating the RL agent in adjusting the strategy for subsequent reasoning accordingly. Please refer to detailed self-evaluation prompts in Appendix K.1.
Action. When addressing difficult and complex problems, humans often employ specific cognitive strategies. For example, we break down complex tasks into smaller components and review previous steps when encountering anomalies in the solution. As evidenced by previous research, understanding and applying these cognitive strategies can significantly enhance the reasoning capabilities of LLMs (Wu et al., 2024; Xue et al., 2024).
With inspiration from human cognition, we design five “basic logic blocks” that can be flexibly combined, constituting the action space in our MDP. By selecting and cascading the blocks, the agent thereby constructs flexible logical structures, paving reasoning pathways from the initial problem to final solutions. In detail, the basic logic blocks include:
  • Reason one step: Perform reasoning for a single step of the current task, which may not directly lead to the final answer but contributes to the overall process.
  • Decompose: Break the current task into simpler subtasks and execute them sequentially. Then, we prompt the LLM to briefly summarize the results of these subtasks as the final result of this action.
  • Debate: Generate multiple plans or approaches for the task at hand and compare them to identify the most promising one. Then, we prompt the LLM to reason one step further based on the selected plan.
  • Refine: Review and revise the current reasoning step to improve clarity and correctness.
  • Terminate: Based on all the previous steps, provide the final answer to the original problem and show it in a specified format. This action marks the conclusion of the reasoning process.
We illustrate the structures and detailed prompts for each blocks in Figure 1 and Appendix K.2.
Reward. To evaluate the quality of the intermediate results after the agent selects an action during the long-sequence reasoning process, we employ the Process Reward Model (PRM) to score the intermediate results and set the PRM score of the intermediate result after each action as the single-step reward for this action.
State Transition. In our MDP design, the state transition is straightforward. During the reasoning process, executing a specific action based on the current problem-solving state is to prompt the LLM to continue reasoning using the logical structure corresponding to that action. After reasoning, the new problem-solving state is obtained through the aforementioned self-evaluation approach.
Also, we impose a few simple restrictions on the state transition, ensuring the correctness and rationality of the constructed logical structures. First, once the answer is already presented in the response after executing some action, no further actions are permitted except for “Terminate”. Second, the “Refine” action is automatically converted to “Reason one step” when it appears as the first action, as no refinement to the original problem is needed. Finally, to avoid the reasoning process being too long, we limit the maximum number of actions, and after reaching the limitation, the “Terminate” action will be automatically executed.
3.3 Training of the Navigator Model
Within the MDP framework outlined above, given a specific LLM type and a kind of reasoning task, we train the navigator model. To enhance the learning for challenging reasoning tasks, we extract hard questions from the training set of the target task, i.e., questions that the LLM cannot answer when directly prompted. Then, we use these problems for training the navigator model, from which we randomly select one in each episode and repeat it multiple times. We employ the Double-Dueling-DQN algorithm (Van Hasselt et al., 2016; Wang et al., 2016), and include the online and target network technique to improve the stability during training (Mnih et al., 2015).
We illustrate the training and inference pipeline of RLoT in Figure 1. During training, we hire PRM to provide reward signals. The parameters of both the PRM and the LLM are kept fixed from pre-trained models, and only parameters of the navigator model, i.e., the RL agent, are updated. This significantly reduces the computational cost, making the training process highly efficient. After training, the PRM model is no longer required, and the trained navigator model is used directly. Given an intermediate reasoning state, the navigator selects an action, which is then used to prompt the LLM to continue reasoning using the logical structure associated with the selected action. By repeating this, the navigator model is able to guide the LLM in solving challenging reasoning tasks with task-specific logical structures.
4 Experiments
4.1 Experimental Settings
Reasoning Tasks. We conduct a comprehensive evaluation of our RLoT method across a wide range of reasoning tasks, encompassing benchmarks in mathematics, STEM, and commonsense question answering. For the mathematics domain, we adopt Olympic-level datasets, the AIME24  [^1] https://huggingface.co/datasets/AI-MO/aimo-validation-aime and AMC23  [^2] https://huggingface.co/datasets/AI-MO/aimo-validation-amc, as well as elementary math datasets GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021c). These are widely recognized as representative benchmarks. In the STEM tasks, we test on the MMLU-STEM (Hendrycks et al., 2021b; a) and GPQA (Rein et al., 2023) datasets, which span various STEM domains and a range of difficulty levels. To evaluate the commonsense reasoning ability, we employ the StrategyQA (Geva et al., 2021) benchmark, which presents challenging multi-hop questions across diverse contexts. These benchmarks cover various domains, difficulties, and task types, forming a systematic evaluation of the reasoning ability.
LLMs. Our RLoT framework is designed to be independent of specific LLMs, allowing it to be compatible with any off-the-shelf LLM. To evaluate this, we test our approach using four representative LLMs: Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct (Yang et al., 2024a), Llama3.1-8B-Instruct (Dubey et al., 2024), GPT-4o-mini (Hurst et al., 2024), and DeepSeek-R1-Distill-Qwen-7B (DeepSeek-AI et al., 2025). In this paper, we mainly focus on sub-10B LLMs, which are often constrained in handling complex reasoning tasks due to their relatively smaller size. We expect that our RLoT design can substantially enhance smaller LLMs by adaptively generating task-specific logical structures at inference time, thereby making their reasoning capabilities comparable to, or even exceeding, those of much larger LLMs. Without the need to modify the LLM’s parameters, this approach will be fairly computationally efficient. Note that we abandon the results of Llama3.1-8B-Instruct on the Olympic-level datasets since the capability of this base LLM is too limited to solve these challenging problems.
Baselines. We compare RLoT against various baselines designed to enhance LLM reasoning at inference time. First, we evaluate single-round question-answering techniques, including direct question answering (Direct QA), zero-shot Chain-of-Thought (Zero-shot CoT), and few-shot Chain-of-Thought (Few-shot CoT) (Wei et al., 2022). For Zero-shot CoT, we employ prompts with “Let’s think step by step”, and for Few-shot CoT, we include specific few-shot examples for each benchmark in the prompts as outlined in prior work (Yang et al., 2024b; Fu et al., 2023; Rein et al., 2023; Wei et al., 2022). Additionally, we consider multi-round techniques, including self-consistent Chain-of-Thought (CoT-SC) (Wang et al., 2023) and Tree-of-Thoughts (ToT) (Yao et al., 2023). Following the original settings, we perform majority voting across four reasoning samples in CoT-SC, and we implement a logical tree with two layers and five nodes per layer in ToT. These multi-round approaches facilitate comparison and voting across diverse reasoning paths, thereby enhancing the reasoning capability of LLMs in complex tasks.
Implementation of RLoT. In our implementation, the navigator model is a simple three-layer multilayer perceptron (MLP) with the Dueling Network architecture (Wang et al., 2016). The model merely contains 2,566 parameters in total, where the lightweight design ensures efficient training and inference. Following existing works like CoT-SC, we perform multiple repeated inferences for each task and filter out the trails that do not meet self-consistency, enhancing the robustness of reasoning.
To obtain reward signals for actions, we utilize the Math-Shepherd as the PRM (Wang et al., 2024c). This model is trained by previous researchers using automatically constructed process-wise supervision data, reducing heavy reliance on manual annotation and thereby achieving remarkable performance. For reproducibility, please refer to more detailed settings and the illustration of the whole pipeline in Appendix J.
Table 2: Overall evaluation of RLoT’s capability to enhance multiple LLMs’ reasoning across different tasks. The bold numbers indicate the best performance in each group of experiments, and the underlined numbers indicate the best baseline method.
LLM Method Olympic math Elementary math STEM Commonsense Average
AIME24 AMC23 MATH GSM8K GPQA MMLU-STEM StrategyQA
Qwen2.5-14B-Instruct Direct QA 13.33 57.50 78.62 93.93 36.60 85.38 72.34 62.53
Zero-shot CoT 16.67 62.50 78.56 94.23 38.39 85.63 75.98 64.57
Few-shot CoT 6.67 55.00 80.00 94.80 45.50 85.06 78.60 63.66
CoT-SC 6.67 47.50 80.04 94.08 45.54 86.71 80.06 62.94
ToT 10.00 55.00 79.50 93.78 45.08 86.55 78.17 64.01
RLoT (ours) 23.33 65.00 80.38 94.16 51.34 88.93 81.22 69.19
Qwen2.5-7B-Instruct Direct QA 10.00 42.50 74.64 91.58 31.25 80.94 68.85 57.11
Zero-shot CoT 16.67 55.00 74.86 91.58 34.15 81.00 71.17 60.63
Few-shot CoT 13.33 45.00 75.50 91.60 36.40 81.16 74.38 59.62
CoT-SC 6.67 50.00 76.36 92.12 38.84 83.25 78.45 60.81
ToT 13.33 50.00 73.80 91.35 36.60 82.62 70.45 59.74
RLoT (ours) 23.33 60.00 76.7 92.87 44.64 85.06 79.04 65.95
Llama3.1-8B-Instruct Direct QA 49.12 84.83 32.14 71.58 70.74 61.68
Zero-shot CoT 49.74 85.37 32.80 72.85 73.07 62.77
Few-shot CoT 48.52 84.50 33.03 70.66 72.05 61.75
CoT-SC 51.74 87.04 33.48 73.29 78.89 64.89
ToT 51.93 86.80 32.24 74.72 71.47 63.43
RLoT (ours) 56.56 90.07 46.88 80.56 84.42 71.70
GPT-4o-mini Direct QA 13.33 55.00 76.58 93.33 43.08 85.70 77.00 61.17
Zero-shot CoT 6.67 67.50 76.76 93.93 40.20 85.76 78.17 64.14
Few-shot CoT 6.67 57.50 75.46 93.48 35.94 85.82 80.06 62.13
CoT-SC 6.67 45.00 76.84 93.63 46.42 86.55 82.53 62.52
ToT 6.67 50.00 74.30 93.33 44.42 85.92 76.42 61.58
RLoT (ours) 20.00 70.00 77.36 93.86 54.02 88.23 82.68 69.45
DeepSeek-R1-
Distill-Qwen-7B
Direct QA 46.67 60.00 92.27 95.74 54.47 83.61 79.48 73.18
Zero-shot CoT 53.33 62.50 91.48 95.20 53.13 85.28 78.89 74.26
Few-shot CoT 56.67 67.50 92.54 94.38 56.47 87.47 79.33 76.34
CoT-SC 56.67 67.50 95.54 96.13 60.94 89.03 82.82 78.38
ToT 50.00 55.00 95.18 94.54 55.13 86.55 80.91 73.90
RLoT (ours) 63.33 77.50 96.56 98.94 67.19 90.77 86.17 82.92
4.2 Overall Performance
We train the navigator model with Qwen2.5-14B-Instruct on the MATH benchmark, where the learning curve shown in Appendix A indicates good convergence. We use the obtained navigator model to enhance the reasoning of multiple LLMs across different reasoning tasks and present the results in Table 2. For the baseline performance, we prioritize using the results reported in the official technical reports of each LLM (Yang et al., 2024a; Dubey et al., 2024; Hurst et al., 2024), if available. Otherwise, we evaluate the performance through our own experiments following the settings established in previous works (Fu et al., 2023; Yang et al., 2024b; Yao et al., 2023).
The results show that our method performs well, consistently outperforming the established inference-time baselines across almost all reasoning tasks when combined with various LLMs. Notably, our approach brings about substantial improvements in the GPQA benchmark, which is a challenging task where LLMs generally perform poorly. Specifically, when implemented with Llama3.1-8B-Instruct, we achieve a 13.4% performance boost. Among the baselines, CoT-SC performs the best across most tasks. Meanwhile, we find that despite being more complex in design, ToT performs poorly on many reasoning tasks, which is also reported in previous studies (Wu et al., 2024; Zhang et al., 2024a; Qi et al., 2024). Furthermore, we report the computational overhead and implementation latency of RLoT in Appendix C, and also compare RLoT with more recently proposed test-time scaling baselines in Appendix E. By directly generate specific logical structures for each question without requiring search-and-trial, our method reached the best performance while maintain a low cost.
4.3 Parameter Size Efficiency
In this section, we demonstrate the parameter size efficiency of our RLoT method in enhancing the reasoning capability of sub-10B LLMs, making them comparable to LLMs with several times more parameters. Specifically, we select three 10B LLMs, including Qwen2.5-14B-Instruct, Llama3.1-8B-Instruct, and GPT-4o-mini, and their respective large-scale counterparts. It is worth noting that the parameter size of GPT-4o series models was estimated in previous studies (Abacha et al., 2024).
[…]
4.4 Transferability
To better illustrate the transferability of our navigator model, we conduct further experiments regarding transferring across different LLMs and reasoning tasks, respectively.
Transfer across Different LLMs. To verify the transferability of RLoT across different LLMs, we respectively train navigator models with three different LLMs, namely Qwen2.5-14B-Instruct, Llama3.1-8B-Instruct, and GPT-4o-mini, on the MATH benchmark. Without any fine-tuning, we cross-test the obtained navigator models to enhance other LLMs on the MATH benchmark and present the results in Table 3.
The results indicate that the trained navigator model exhibits strong transferability across different LLMs. When implementing the navigator model to enhance the reasoning capabilities of a specific LLM, we find that, regardless of whether the navigator model is trained on the same LLM or a different one, its performance remains consistent. In all cases, the enhanced LLM outperforms the best well-known inference-time baseline.
Table 3: Evaluation of RLoT’s transferability across different LLMs. We train navigator models with three different LLMs on the MATH benchmark and cross-test the obtained navigator models with other LLMs. We also list CoT-SC, the best baseline method, for comparison.
Method Train Test
Qwen2.5-14B-Instruct
@MATH
Llama3.1-8B-Instruct
@MATH
GPT-4o-mini
@MATH
RLoT (ours) Qwen2.5-14B-Instruct@MATH 80.38 56.56 77.36
Llama3.1-8B-Instruct@MATH 81.48 53.60 78.14
GPT-4o-mini@MATH 80.84 56.94 78.08
CoT-SC 80.04 51.74 76.84
Transfer across Different Reasoning Tasks. To verify the transferability of RLoT across different reasoning tasks, we respectively train navigator models with Qwen2.5-14B-Instruct on three different benchmarks, namely MATH, GPQA, and StrategyQA. Without any fine-tuning, We cross-test the obtained navigator models to enhance the reasoning capabilities of Qwen2.5-14B-Instruct on the other tasks and present the results in Table 4.
The results indicate that the trained navigator model owns strong transferability across different reasoning tasks. When utilizing the navigator model to enhance the reasoning capabilities of LLMs for a specific task, we observe that, regardless of whether the navigator model is trained on the same task or a different one, its performance remains largely consistent. In most cases, it surpasses the best-performing inference-time baseline.
Furthermore, we find that the navigator models trained on mathematical (MATH) and STEM (GPQA) problems exhibit better transferability to each other. However, the transferability between the navigator trained on commonsense problems (StrategyQA) and the former two is relatively limited, which is intuitive given the inherent relations and differences between domains of mathematics, STEM, and commonsense reasoning.
Table 4: Evaluation of RLoT’s transferability across different tasks. We train navigator models with Qwen2.5-14B-Instruct on three different tasks and cross-test the obtained navigator models on other tasks. We also list CoT-SC, the best baseline method, for comparison.
Method Train Test
Qwen2.5-14B-Instruct
@MATH
Qwen2.5-14B-Instruct
@GPQA
Qwen2.5-14B-Instruct
@StrategyQA
RLoT (ours) Qwen2.5-14B-Instruct@MATH 80.38 51.34 81.22
Qwen2.5-14B-Instruct@GPQA 80.76 53.57 80.64
Qwen2.5-14B-Instruct@StrategyQA 79.94 50.22 81.37
CoT-SC 80.04 45.54 80.06
4.5 Typical Reasoning Patterns
Table 5: Typical patterns in the task-specific logical structures generated by RLoT (“Reason” is short for “Reason one step”).
Task MATH GPQA StrategyQA
2-step pattern Reason-Refine Reason-Refine Reason-Debate
Reason-Decompose Reason-Debate Reason-Refine
3-step pattern Decompose-Refine-Reason Reason-Refine-Debate Reason-Decompose-Debate
Reason-Refine-Debate Reason-Decompose-Refine Reason-Refine-Debate
The experimental results above have demonstrated RLoT’s capability to enhance LLM reasoning with task-specific logical structures. In Table 5, we summarize typical patterns observed in the logical structures generated by RLoT when solving tasks from the MATH, GPQA, and StrategyQA benchmarks using Qwen2.5-14B-Instruct.
From the 2-step patterns, we observe that the Reason-Refine mode emerges frequently. Particularly in MATH and GPQA, which require massive mathematical calculations, this pattern compensates for the LLMs’ relatively poor calculation abilities, leading to more reliable results. In the 3-step patterns, operations like Decompose and Debate are frequently employed, which help break down challenging problems or facilitate discussions to explore potential solutions. Additionally, the Refine operation is often used before and after the Decompose and Debate steps, ensuring correct integration with preceding and succeeding reasoning processes. These typical reasoning patterns exhibit strong interpretability, further validating the capability of RLoT to generate task-specific logical structures that enhance LLM reasoning. For better understanding, we provide an example in Appendix B to illustrate how these logical structures empower the correct solution to a problem.
4.6 Ablation Studies, Analyses, and Discussions
To better justify our designs, we conducted a series of ablation studies, analyses, and discussions. We discuss the reliability of the self-evaluation state design in Appendix F. We analyze the necessity of employing RL to train the navigator model in Appendix G. We illustrate the role of the process reward model (PRM) in training the navigator in Appendix H. Also, we verify the effectiveness of our logic block designs and quantify the contribution of each logic block in Appendix I.
5 Related Works
5.1 Inference-time reasoning of LLMs
To improve LLMs’ reasoning capability, plentiful research has investigated inference-time techniques without the need for model updates. On the one hand, predefined external logical structures are widely applied as basic solutions. Most notably, Chain-of-Thought (CoT) (Wei et al., 2022) incorporates intermediate reasoning steps within the prompt, enhancing the model’s abilities in complex tasks. As a subsequent advancement, CoT with Self-Consistency (CoT-SC) (Wang et al., 2023) further refines CoT by generating multiple reasoning chains and filtering out those that do not meet self-consistency, thus increasing the reliability. Moreover, the concept of Tree-of-Thoughts (ToT) (Yao et al., 2023) and Graph-of-Thoughts (GoT) (Besta et al., 2024) has been introduced, where the reasoning process is represented as a graph, enabling exploration and backtracking from more promising outcomes. Despite their success, these methods rely on task-agnostic logical structures that are applied uniformly across diverse tasks, lacking flexibility.
On the other hand, recent researchers have proposed more adaptive inference-time approaches. From hiring the decompose-analyze-rethink procedure (Xue et al., 2024) to utilizing Monte Carlo Tree Search to discover more effective logical structures (Wu et al., 2024; Wang et al., 2024a), these approaches dynamically compose appropriate logical structures for specific tasks to enhance the reasoning performance. However, the search-and-trail design incurs massive extra computational cost, limiting the efficiency of reasoning. In contrast, the proposed RLoT method employs RL to train a navigator agent. With the trained navigator, our design can directly select and combine basic logic blocks and generate task-specific logical structures, enabling more adaptive reasoning while eliminating the searching cost.
5.2 RL and LLMs
RL has become vital in the development of LLMs (Xu et al., 2025; Hao et al., 2025). One important direction is aligning LLMs with human preferences, where the key method is Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022). In RLHF, LLMs are fine-tuned as actors in RL based on feedback derived from human preferences. Recently, extensive helpful and harmless LLMs have been created using RLHF (Bai et al., 2022; Casper et al., 2023). Beyond this, RL is also applied to enhance the reasoning capability of LLMs. In this context, reward signals are derived from Outcome Reward Models (ORM) (Kazemnejad et al., 2024) or Process Reward Models (PRM) (Lightman et al., 2023; Wang et al., 2024c; Luo et al., 2024), providing feedbacks for the LLMs’ reasoning process during fine-tuning. By applying RL to LLMs with these rewards, LLMs can iteratively improve their performance in multi-step reasoning tasks (Havrilla et al., 2024a; b; Shao et al., 2024; DeepSeek-AI et al., 2025).
In summary, existing RL techniques have significantly enhanced LLMs’ capabilities by updating model parameters, yet such fine-tuning demands substantial computational resources when applied to each pre-trained LLM. In this paper, the proposed RLoT method applies RL at inference time, leveraging the power of RL to train a lightweight navigator rather than the entire parameters of LLMs, our method achieves low cost and wide compatibility with various pre-trained LLMs.
6 Conclusions
In this paper, we propose RL-of-Thoughts (RLoT), an inference-time technique that utilizes RL to train a navigator model, which adaptively constructs task-specific logical structures, and thereby enhances the reasoning capabilities of LLMs. Through extensive experiments across various benchmarks, we demonstrate the effectiveness of our method in improving the reasoning capability of various widely known LLMs. Additionally, we show the strong transferability and efficiency of our approach, where the trained navigator model can effectively transfer across different LLMs and unseen reasoning tasks. With fewer than 3K parameters, our navigator model enables multiple sub-10B LLMs to attain performance comparable to larger LLMs with up to 10 times the parameter size. Our work highlights the potential of RL at inference time in enhancing the reasoning capabilities of LLMs, paving the way for more adaptive and efficient LLM reasoning in the future.
Ethics statement
This study uses fully open-source or publicly available models and datasets, adhering to their respective licenses. All resources are properly cited in Section 1 and Section 4. The selected datasets and models are well-established, representative, and free from bias or discrimination.
Reproducibility statement
For Reproducibility, we describe the general experimental settings in Section 4; we list the implementation details in Appendix J; and our source code are anonymously open source at https://anonymous.4open.science/r/RL-LLM-Reasoning-1A30.
Appendix A Learning Curves
In Figure 2, we show the learning curves during RL training of all navigator models used in experiments in the main text. We use a sliding window averaging to smooth the reward, and from the learning curves, we observe good convergence of the RL training.
Figure 2: Learning curves during RL training of all navigator models.
Appendix B Case Study
Figure 3: A case study comparing few-shot CoT and RLoT on a representative problem in the MATH dataset. Results are both generated by Qwen2.5-14B-Instruct.
As shown in Figure 3, we analyze a representative case to illustrate the detailed working procedure of RLoT. The problem is sampled from the test set of MATH (Hendrycks et al., 2021c) benchmark, which involves calculating the area between two functional curves, requiring a combination of algebraic and geometric knowledge. While the Few-shot CoT method attempts to generate the answer in one step, RLoT adopts a task-specific reasoning pathway, which includes Reason-Refine-Debate, a typical pattern for solving complex math problems shown in Section 4.5.
Both Few-shot CoT and RLoT successfully calculate the intersection points of the curves. However, the response of Few-shot CoT fails to accurately identify the shape of the shadowed area, leading to incorrect bases for the triangles and, ultimately, an incorrect result. In contrast, RLoT utilizes the Refine action to clarify the intermediate results and correctly identifies the shadowed area as a rhombus. Subsequently, the Debate action proposes multiple plans and prioritizes directly using the area formula for a rhombus over dividing it into triangles. By avoiding unnecessary geometric division, RLoT achieves the correct answer.
Appendix C Computational Overhead and Latency
As shown in Table 6, RLoT improves the reasoning ability of LLMs with a low cost comparing with complex inference-time techniques. The results show that at inference-time, RLoT’s token consumption only accounts for 78 % of ToT, and the consumption is comparable to CoT-SC. This is because the trained navigator can directly select appropriate logic blocks without searching like ToT, improving the performance and reducing cost. Besides, our navigator with <3k parameters has almost negligible inference cost itself.
Though RL training incurs extra costs, our lightweight navigator makes training efficient, and the trained navigator can transfer across tasks, diluting the cost. Specifically, our navigator is trained for 3000 episodes, and each episode consumes an average of 5004 input and 1448 output tokens. Shared across all test questions, it only adds less than 5% input and output tokens per question.
Table 6: Token consumption per question across tasks. The bold numbers indicate our method.
Model Method AIME24 AMC23 MATH GSM8K GPQA MMLU-STEM StrategyQA Average
Input Output Input Output Input Output Input Output Input Output Input Output Input Output Input Output
Qwen2.5-14B-Instruct Direct QA 653 1433 619 1131 46 362 55 183 159 460 82 210 30 259 235 577
Zero-shot CoT 667 1489 633 1007 58 367 66 206 175 546 87 334 33 376 246 618
Few-shot CoT 1955 1341 1921 940 466 336 1089 135 1557 589 1114 259 421 223 1218 546
CoT-SC 7823 5330 7685 4109 2014 1320 4339 543 6295 1945 4448 1039 1586 897 4884 2169
ToT 5994 10844 3836 8029 4983 6277 3570 4312 6063 7404 4797 5363 5762 6491 5001 6960
RLoT (ours) 22004 3888 17963 3463 3735 1109 2310 634 5501 1485 2923 791 2348 756 8112 1732
Qwen2.5-7B-Instruct Direct QA 642 1714 644 1443 44 366 58 182 159 546 66 404 22 279 234 705
Zero-shot CoT 667 1385 663 1180 47 393 78 222 184 522 82 413 32 401 250 645
Few-shot CoT 4230 1789 3875 1501 483 581 1074 188 1556 442 1128 221 396 348 1820 724
CoT-SC 16890 7131 15341 5821 1932 1443 4379 750 6290 1800 4433 883 1612 1391 7268 2746
ToT 5289 9180 3741 7621 6679 7525 5792 5301 5771 6489 5058 4669 3807 2013 5162 6114
RLoT (ours) 23093 4168 12876 2896 1437 807 1283 476 2112 889 1028 434 1735 606 6223 1468
Llama3.1-8B-Instruct Direct QA 44 437 53 180 174 811 71 406 13 201 71 407
Zero-shot CoT 73 495 64 215 175 661 91 365 20 418 85 431
Few-shot CoT 510 472 1101 173 1568 769 1127 168 417 272 945 371
CoT-SC 2049 1885 4325 808 6120 3078 4429 808 1582 1071 3701 1530
ToT 3474 4307 5212 4780 5615 3794 5331 5999 4320 4868 4790 4750
RLoT (ours) 4481 1981 2842 1028 9818 3387 4408 1486 3049 1094 4920 1795
GPT-4o-mini Direct QA 601 1356 571 1184 43 404 54 223 165 397 86 175 13 201 219 563
Zero-shot CoT 614 1435 585 1192 50 422 68 260 160 525 89 326 16 382 226 649
Few-shot CoT 1853 1365 1823 1202 513 408 1107 205 1575 505 1119 275 409 254 1200 602
CoT-SC 7414 5493 7294 5097 2030 1554 4442 823 6305 2027 4360 1005 1601 1014 4778 2430
ToT 4639 4929 5726 4521 4534 5649 5890 5810 5175 5841 4950 5324 3833 1467 4964 4792
RLoT (ours) 26529 4778 19513 3973 3483 1575 3069 1048 4372 1550 1882 710 1612 673 8637 2044
DeepSeek-R1-
Distill-Qwen-7B
Direct QA 879 11793 597 4686 541 4960 533 1364 808 5462 521 2094 436 1268 616 4518
Zero-shot CoT 906 12129 640 5140 565 4697 547 1381 822 5443 574 2012 450 1344 643 4592
Few-shot CoT 2360 7778 2086 5120 1938 3301 3253 1303 4390 5194 2063 1890 1197 1275 2470 3694
CoT-SC 9240 32005 8556 21859 7749 13071 12976 5123 17562 20553 7883 7820 4791 5094 9822 15075
ToT 5879 34616 4474 45524 3076 25913 6680 26033 4481 19222 2158 9604 8393 11697 5020 24658
RLoT (ours) 13556 21870 15684 31915 4003 8450 8347 12389 3734 18787 6732 8613 6445 10546 8357 16081
Average Direct QA 694 4074 608 2111 144 1306 151 426 293 1535 165 658 103 442 308 1507
Zero-shot CoT 714 4110 630 2130 159 1275 165 457 303 1539 185 690 110 584 324 1541
Few-shot CoT 2600 3068 2426 2191 782 1020 1525 401 2129 1500 1310 563 568 474 1620 1317
CoT-SC 10342 12490 9719 9222 3155 3855 6092 1609 8514 5881 5111 2311 2234 1893 6452 5323
ToT 5450 14892 4444 16424 4549 9934 5429 9247 5421 8550 4459 6192 5223 5307 4996 10078
RLoT (ours) 21296 8676 16509 10562 3428 2784 3570 3115 5107 5220 3395 2407 3038 2735 8049 5071
RLoT plus training cost 21649 8779 16863 10664 3782 2887 3924 3217 5461 5322 3749 2509 3392 2837 8403 5174
Besides token consumption, we show the average solving time (min) per problem along with the standard deviation across 3 repeats in Table 7.
Table 7: Average solving time (minutes) per problem across various benchmarks. Standard deviations from 3 repeats are shown in parentheses.
Method AIME24 AMC23 MATH GSM8K GPQA MMLU-STEM StrategyQA Average
Direct QA 0.44(0.02) 0.25(0.01) 0.13(0.00) 0.05(0.00) 0.17(0.01) 0.08(0.00) 0.05(0.00) 0.17(0.01)
Zero-shot CoT 0.44(0.02) 0.25(0.02) 0.13(0.01) 0.06(0.00) 0.17(0.02) 0.08(0.00) 0.06(0.00) 0.17(0.01)
Few-shot CoT 0.52(0.03) 0.42(0.02) 0.17(0.01) 0.18(0.01) 0.33(0.02) 0.17(0.02) 0.10(0.01) 0.27(0.02)
CoT-SC 2.09(0.13) 1.74(0.12) 0.64(0.03) 0.71(0.07) 1.32(0.09) 0.68(0.05) 0.38(0.02) 1.08(0.07)
ToT 1.86(0.07) 1.91(0.15) 1.33(0.12) 1.35(0.11) 1.28(0.06) 0.98(0.10) 0.97(0.11) 1.38(0.10)
RLoT (ours) 2.75(0.21) 2.48(0.22) 0.57(0.02) 0.61(0.05) 0.95(0.05) 0.53(0.03) 0.53(0.04) 1.20(0.09)
Our navigator can directly generate a task-specific logical trajectory, reducing the LLM interaction cost. Also, our lightweight navigator itself brings almost negligible cost.
Appendix D Performance Comparison with Larger LLMs
Table 8: Performance comparison between sub-10B LLMs enhanced by RLoT, and larger LLMs with several times of parameters.
LLM Size Method MATH GSM8K GPQA MMLU-STEM StrategyQA Average Gap
Qwen2.5-Instruct 14B Few-shot CoT 80.00 94.80 45.50 85.06 78.60 76.79 -4.13
RLoT (ours) 80.38 94.16 51.34 88.93 81.22 79.21 -1.71
72B Few-shot CoT 83.10 95.80 49.00 89.80 86.90 80.92
Llama3.1-Instruct 8B Few-shot CoT 48.52 84.50 33.03 70.66 72.05 61.75 -12.51
RLoT (ours) 56.56 90.07 46.88 80.56 84.42 71.70 -2.56
70B Few-shot CoT 68.00 95.10 46.70 84.81 76.71 74.26
GPT-4o mini (8B) Few-shot CoT 75.46 93.48 35.94 85.82 80.06 74.15 -4.44
RLoT (ours) 77.36 93.86 54.02 88.23 82.68 79.23 +0.64
(200B) Few-shot CoT 76.60 93.73 53.60 87.90 81.10 78.59
[…]
Appendix E Extended Baseline Comparisons
E.1 Test-time Scaling Baselines
In addition to math reasoning models, we also compare RLoT with multiple models which cover different categories and designs. The details of baselines are listed below.
Fixed reasoning patterns. Baselines from this category adopts a fixed workflow to obtain the answer, which
  • DeAR (Xue et al., 2024): The problem is consequently decomposed, analyzed and refined in this workflow. The workflow follows a recurrent manner to a detailed decomposition of original problem.
  • Self-Refine (Madaan et al., 2023): This method adds a simple refine step after every reasoning step to correct potential mistakes.
Tree search.Tree-based models divide the reasoning task into sub-steps, and searches for a best path to the final answer. Tree-based methods usually requires multiple rounds Q/A, resulting in high reasoning cost.
  • One-step-greedy: Using PRM model to greedily select the best thought at each step.
  • Basic Monto-carlo Tree Search: An effective tree-searched method that explore the tree via multiple roll-outs. The MCTS is also supervised by the PRM.
  • Litesearch (Wang et al., 2024a): An low-cost tree search method that assigns larger search budget to a more promising nodes.
  • $\mathbf{Q}^{*}$ (Wang et al., 2024b): $\mathbf{Q}^{*}$ adopts a value function that estimate the "distance" between the current thought and the correct answer. Then an $\mathbf{A}^{*}$ algorithm, which is widely applied in shortest path problem, is used to finish tree-search.
  • rStar (Qi et al., 2024): Use Monte Carlo Tree Search during inference to select actions.
  • AFlow (Zhang et al., 2024b): An automated framework that uses MCTS to efficiently explore LLM agentic workflow.
Others.
  • DSPy (Khattab et al., 2023): A programming model that can express and optimize sophisticated LM pipelines.
  • Graph of Thoughts (GoT) (Besta et al., 2024): A graph-based model which allows arbitrary combination of thoughts according to the manually designed graph architecture. We carefully design graphs for each specific dataset.
  • Buffer of Thoughts (BoT) (Yang et al., 2024d): A Retrieval-Augmented Generation (RAG) model that enhance reasoning via reasoning patterns in the buffer. These patterns are continuously updated during inference time.
Table 9: Extended baseline comparisons. The bold numbers indicate the best performance in each group of experiments, and the underlined numbers indicate the best baseline method.
LLM Method GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
DeAR (Xue et al., 2024) 88.38 36.16 71.13 65.22
Self-Refine (Madaan et al., 2023) 88.55 32.59 73.94 65.03
Tree search One-step-greedy 89.08 32.59 73.07 64.91
Basic MCTS 90.07 34.82 75.69 66.86
Litesearch (Wang et al., 2024a) 89.76 33.04 75.84 66.21
Q* (Wang et al., 2024b) 91.51 34.60 77.00 67.70
rStar (Qi et al., 2024) 91.96 38.62 77.15 69.24
AFlow (Zhang et al., 2024b) 91.36 37.72 76.56 68.55
DSPy) (Khattab et al., 2023) 88.32 34.15 74.09 65.52
Graph of Thoughts (GoT) (Besta et al., 2024) 88.96 33.04 75.42 65.81
Buffer of Thoughts (BoT) (Yang et al., 2024d) 92.35 34.51 74.33 67.06
RLoT (ours) 92.87 44.64 79.04 72.18
As shown in Table 9,we compare these methods on representative datasets, GSM8K, GPQA, and StrategyQA, which covers three different domains. Results show that our model outperforms all baseline methods among all datasets.
By flexibly organizing logic blocks, our design can outperform these baselines. To compare the efficiency and effectiveness, we show the token consumption per question for representative baselines:
Table 10: Token consumption and performance scores per question for representative baselines.
LLM Method GSM8K GPQA StrategyQA Average
Score Token Score Token Score Token Score Token
Qwen2.5-Instruct-7B DirectQA 91.58 240 31.25 705 68.85 301 63.89 415
Self-Refine 88.55 678 32.59 1784 73.94 939 65.03 1134
rStar 91.96 2636 38.62 4127 77.15 2987 69.24 3483
RLoT (ours) 92.87 1759 44.64 3001 79.04 2341 72.18 2367
Unlike search-based methods that require multiple explorations, our design directly generates task-specific logical trajectories, allowing for better performance with a similar or lower cost.
E.2 Math Domain Baselines
To better demonstrate the effectiveness of our design, we compare RLoT with more baselines that have been specifically designed for math reasoning tasks. We involve fine-tuning designs which are computationally expensive.
  • AceMath (Liu et al., 2024): A supervised fine-tuning model designed specifically for mathematical reasoning. It first develops a math-specialized reward model using public datasets and then performs fine-tuning and reasoning guided by this reward model.
  • PFPO (Jiao et al., 2024): A supervised fine-tuning approach guided by pseudo reward feedback. The feedback is generated either through a self-consistency mechanism or with the assistance of more powerful LLMs.
We also include inference-time designs that hire complicated algorithms like Monte-Carlo Tree Search (MCTS).
  • LLM2 (Yang et al., 2024c): A lightweight model aimed at enhancing LLM reasoning during inference. It achieves this by training a verifier to distinguish and prioritize better LLM-generated responses.
  • LLaMA-Berry (Zhang et al., 2024a): An inference-time Monte Carlo Tree Search (MCTS) method that explores reasoning paths using a trained reward model.
  • HiAR-ICL (Wu et al., 2024): A method that matches problems with multiple reasoning templates at inference time. These templates are previously generated using MCTS on a subset of the dataset.
Table 11: Performance comparison of RLoT’s with more baselines. The bold numbers indicate the best performance in each category.
Category Method MATH GSM8K Average
Fine-tuning AceMath (Liu et al., 2024) 64.42 90.45 77.44
PFPO (Jiao et al., 2024) 57.80 89.60 73.70
Inference-time LLM2 (Yang et al., 2024c) 48.60 88.00 68.30
LLaMA-Berry (Zhang et al., 2024a) 54.80 89.80 72.30
HiAR-ICL (Wu et al., 2024) 55.00 90.70 72.85
RLoT (ours) 56.56 90.07 73.32
We test with Llama3.1-8B-Insturct on MATH and GSM8K benchmarks and show the results in Table 11. For the baselines, we use the performance reported in the original papers, and for RLoT, we test with the same navigator model as in Section 4.2 in the main text. From the results, we can observe that our method outperforms all inference-time baselines. Meanwhile, with substantially lower computational consumption, it reaches comparable performance to some fine-tuning methods that require modification on the parameters of LLMs.
Appendix F Analyses on the Reliability of Self-evaluation State
Previous work shows that LLMs can evaluate and correct their own outputs (Han et al., 2024). Meanwhile, multiple works (Madaan et al., 2023; Xue et al., 2024) have reached success with self-correcting. Based on the literature, we design the structured self-evaluation that prompts the LLM across granular aspects (Table 1), including the correctness of modeling and calculation. This design aims to be more specific and reliable than the single judgment.
To validate the robustness, we randomly sample 100 intermediate reasoning texts and manually check the self-evaluation. We find that the LLM’s assessment was accurate in 82 out of 100 cases. Also, we deliberately introduced modeling and calculation errors in a reasoning path, and test whether our self-evaluation mechanism is able to correctly identify. As shown in Table 12, the problem is from GSM8K (Cobbe et al., 2021), which includes the relationship of three variables. In addition to the correct reasoning step, we modify it to obtain two wrong reasoning steps, which respectively makes a mistake on modeling and calculation. The result shows that the LLM is able to identify and classify different kinds of mistakes, demonstrating the effectiveness of self-evaluation state extraction.
Table 12: An example of self-assessing of intermediate states.
Question Reasoning step Self-evaluated correctness
Modeling Calculation
Mark has a garden with
flowers. He planted plants
of three different colors in it.
Ten of them are yellow,
and there are 80% more of
those in purple. There are only
25% as many green flowers as
there are yellow and purple
flowers. How many flowers
does Mark have in his garden?
In this step, we aim at calculate the number
of purple flowers, which is 80% more
than the yellow ones.
We calculate by 10 $\times$ (1 + 0.8) = 18
True True
In this step, we aim at calculate the number
of purple flowers, which is 80% more
than the yellow ones.
We calculate by 10 $\times$ 0.8 = 8
False True
In this step, we aim at calculate the number
of purple flowers, which is 80% more
than the yellow ones.
We calculate by 10 $\times$ (1 + 0.8) = 17
True False
Based on the validation, we consider our structured self-evaluation provides a sufficiently reliable state representation for training the navigator.
Furthermore, we explore alternative designs of self-evaluation:
  • MLP. Use the self-evaluation vector from a fixed LLM as input, only train an MLP as the navigator.
  • LLM (fixed). Directly prompt the LLM to select the action based on the raw reasoning trajectory.
  • LLM (finetune). Use the raw reasoning trajectory as input, train an LLM backbone with an MLP classification head as the navigator.
Table 13: Comparison of alternative self-evaluation designs.
LLM Navigator GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
LLM (fixed) 89.16 37.95 74.38 67.16
LLM (finetune) 88.32 34.15 70.01 64.16
MLP (ours) 92.87 44.64 79.04 72.18
As the results in Table 13 show, LLM (fixed) performs badly since pre-trained LLMs are not specified for the task. LLM (finetune) even performs worse. Our method injects valuable human knowledge for explicit, structured self-evaluation. In contrast, LLM (finetune) requires an implicit evaluation from raw text, which is much more complex and has more parameters to be tuned. This requires substantially more consumption and specialized design to overcome the complexity and uncertainty. Therefore, our method is a comprehensive design at a low cost. While some self-evaluation cases may be inaccurate, the overall improvement demonstrates the general benefit.
Appendix G Analyses on the Role of RL in Training the Navigator
To illustrate the necessity of using RL in training the navigator, we compare the RL navigator with several other decision methods for deciding the next reasoning block. The methods includes the
  • Fixed logic sequence. Fixed "Decompose-Reason-Refine" sequence like human thinking.
  • Supervised-trained navigator. We first collected 20K state-action-score samples with random actions. Then, we trained a model to predict the outcome for each action given a state, and select the action with the highest predicted score during inference.
  • LLM as navigator. Directly prompt another LLM to select the action based on the reasoning context.
We also add:
  • Repeated strong blocks. Repeat Debate or Refine until reaching the answer.
Table 14: Comparison with alternative navigator methods on various benchmarks.
LLM Method GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 66.62
Repeated refine 87.72 32.37 72.34 66.65
Repeated debate 89.99 33.04 72.05 67.51
Fixed logic sequence 88.38 36.16 71.13 67.47
Supervise-trained navigator 89.84 36.83 70.31 67.94
LLM as navigator 89.16 37.95 74.38 69.17
RL-trained navigator (ours) 92.87 44.64 79.04 73.34
The results in Table 14 show that these methods are worse than RLoT:
  • Fixed logic sequence. Rigid, unable to design flexible logical structures for specific tasks.
  • Supervise-trained navigator. The randomly sampled state-action pairs lack the ability of RL to purposely exploit effective parts of the sample space.
  • LLM as navigator. Pre-trained LLMs are not specified for the logic selection task.
These show the importance of using RL for our navigator to flexibly select logic blocks.
Appendix H Analyses on the Role of Process Reward Model (PRM)
As literature (Zhang et al., 2025) indicates that math-trained PRMs perform well on MMLU-STEM, and the performance of PRMs in math and STEM is generally correlated. Our navigator, which is trained with a math PRM, achieves strong performance on GPQA (STEM). This also reflects the generalization capability of PRM across domains. We also test an ORM to train the navigator.
Table 15: Comparison of RLoT using an Outcome Reward Model (ORM) vs. a Process Reward Model (PRM).
LLM Method GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 66.62
RLoT with ORM 91.96 41.52 73.94 70.71
RLoT with PRM (ours) 92.87 44.64 79.04 73.34
The results in Table 15 show that the PRM outperforms the ORM, indicating that the step reward signal from PRM is crucial.
Further, to explore the impact of PRM quality, we test:
  • ORM: As a minimal quality PRM.
  • Degraded PRM: Add a std=0.1 Gaussian noise to our raw PRM.
  • Better PRM: Qwen2.5-Math-PRM-7B, a stronger PRM.
Table 16: Impact of PRM quality on navigator performance.
LLM PRM GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B None-DirectQA 91.58 31.25 68.85 63.89
ORM 91.96 41.52 73.94 69.14
MathShepherd-Mistral-7B + disturb 92.49 42.41 77.15 70.68
MathShepherd-Mistral-7B 92.87 44.64 79.04 72.18
Qwen2.5-Math-PRM-7B 93.10 45.31 80.06 72.82
These results in Table 16 show that a higher-quality PRM is beneficial. However, the navigator is robustly beneficial with a lower-quality PRM, as long as the reward signal is directionally meaningful.
Appendix I Analyses on the Design of Logic Blocks
We first check the test-time scaling for the logic blocks. Results in Figure 4 show that a longer sequence of logic blocks brings a performance gain.
Figure 4: Test-time scaling with different number of logic blocks.
To examine the impact of each block, we remove each of them and train a navigator:
Table 17: Ablation study results for each logic block.
LLM Ablation MATH GPQA StrategyQA Average
Qwen2.5-Instruct-7B w/o Decompose 75.42 31.92 77.00 61.45
w/o Debate 74.02 36.61 77.58 62.74
w/o Refine 75.76 41.29 72.93 63.33
Full RLoT 76.70 44.64 79.04 66.79
GPT-4o-mini w/o Decompose 76.08 41.74 79.33 65.72
w/o Debate 74.68 45.31 78.75 66.25
w/o Refine 76.44 51.56 74.38 67.46
Full RLoT 77.36 54.02 82.68 71.35
The ablation results in Table 17 confirm that all blocks are effective.
We also analyze the impact of different logic blocks on different datasets. Results in Figure 5 show the average PRM reward gain of each logic block. The most useful block varies in different tasks, specifically Debate for math tasks, Decompose for STEM tasks, and refine for common-sense reasoning.
Figure 5: Contribution of each logic block in the reasoning sequence.
Further, we test some possible new blocks:
  • Follow-up-question: Propose and answer a sub-question based on the current reasoning.
  • Rephrasing: Reorganize the current reasoning.
  • Compression: Condense the current reasoning.
Table 18: Performance with additional logic blocks.
LLM Logic blocks MATH GPQA StrategyQA Average
Qwen2.5-Instruct-7B Original in RLoT 76.70 44.64 79.04 66.79
Original + Follow-up-question 76.78 43.97 79.24 66.66
Original + Rephrasing 77.30 45.54 80.20 67.68
Original + Compression 76.06 43.53 77.87 65.82
The results in Table 18 showed that "Rephrasing" is beneficial, while "Follow-up-question" (perhaps self-ask-and-answer has limited effect) and "Compression" (likely due to information loss) are not. This shows the extensibility of RLoT: it provides a framework for the community to design more reasoning blocks to enhance the performance.
Appendix J Implementation Details
J.1 Settings
In this section, we provide the main implementation settings for reproducibility in Table 19. Please refer to our source code at https://anonymous.4open.science/r/RL-LLM-Reasoning-1A30 for the exact usage of each hyper-parameters and more details.
Table 19: Implementation details.
Module Element Detail
System OS Ubuntu 22.04.2
CUDA 11.7
Python 3.11.4
Double-Dueling DQN $\gamma$ 0.9
Number of episodes 3000
Batch Size 64
Interval of target network updating 50
Optimizer Adam
Learning rate 0.01
Learning rate decay 0.5 per 1000 episodes
Replay buffer size 500
Start epsilon 1
Min epsilon 0.0
Epsilon decay 0.9995 per step
RLoT framework Maximum number of actions 5
Number of trails for self-consistency 3
Temperature for LLMs 1.0
J.2 Illustration of the full pipeline
Here we illustrate how every parts in our design work together as an whole pipeline.
1 Data construction
We pick out "hard problems", namely those that the LLM wrongly answers when directly prompted, for training the navigator. Here are more details: we used Qwen2.5-14B-Instruct as a reference model to select "hard problems" from the training sets. The direct prompts are:
Prompts:
{Problem} Please generate the answer for the problem.
(MATH) Wrap the answer with boxed{{answer}}.
(GPQA) End the answer with ’The answer is (CHOICE)’.
(StrategyQA) End the answer with ’YES/NO’.
The statistics on benchmarks involved in our training are shown in Table 20.
Table 20: Statistics on benchmarks involved in our training.
Dataset Total Hard Proportion (%)
MATH 7500 1736 23.15
GPQA 448 255 56.92
StrategyQA 1600 341 21.31
By selecting "hard problems", our navigator dedicates to strategies that enhance LLMs on problems that they cannot solve directly. For our RL approach, we do not construct (input, expected_output) pairs like supervised learning. The training data consists of individual questions. During training, the model is fed with questions, and the outputs are evaluated to generate rewards, which are then fed back to train the model. We will detail the key components and entire process of training below.
2 State self-evaluation (state space)
We obtain the 7-aspect state vector through a prompted self-evaluation mechanism. We feed the problem and the existing reasoning into the LLM with the prompt in Appendix K.1. This design compresses the textual reasoning trajectory into a low-dimensional state vector, making the policy learning process more efficient and tractable.
Here is an input-output example of state self-evaluation, as seen in our case study in Appendix F:
Input
(Problem) Mark has a garden with flowers. He planted plants of three different
colors in it. Ten of them are yellow, and there are 80% more of those in
purple. There are only 25% as many green flowers as there are yellow and
purple flowers. How many flowers does Mark have in his garden?
(Existing reasoning) In this step, we aim at calculating the number of purple
flowers, which is 80% more than the yellow ones. We calculate by 10 * (1 + 0.8) = 17
Output
(State vector) {Modelling: True, Calculation: False, ...}
3 Logical blocks (action space)
First, these logic blocks are inspired by well-established cognitive strategies in human problem solving. "Decompose," "Debate," and "Refine" are the core actions, supplemented by "Reason one step" and "Terminate" to ensure complete reasoning flows.
For instance, "Decompose" reflects the widely studied divide-and-conquer approach, while "Debate" aligns with strategies involving comparative evaluation (Wang & Chiew, 2010). Also, they are empirically effective. Prior studies (Madaan et al., 2023; Xue et al., 2024) have demonstrated that "Decompose" and "Refine" can significantly enhance the performance of LLMs on complex tasks.
Second, while preserving the high-level cognitive intuition, we have operationalized them into specific primitives that the LLM can execute. The key differences are:
  • Formalization with prompts: Each block corresponds to a prompt template that instructs the LLM to perform a specific function (Appendix K.2).
  • Standard Input/Output: All blocks operate on a unified interface. The input is the original problem and the existing reasoning, and the output is a new segment of reasoning. This allows flexible composition of logic blocks.
Third, here is an input-output example for the ’Debate’ block (as Appendix B):
Input
(Problem) How many square units are in the region satisfying the inequalities
\$y $\ge$ |x|\$ and \$y $\le$ -|x|+3\$?
(Existing reasoning) Let’s start by analyzing the given inequalities. The next
step is to determine the points of intersection... Now we know the points of
intersection are...
Output
(New reasoning) The most promising plan is: Given the diagonals of the rhombus,
we can now calculate the area.
The contributions of the logic blocks are discussed in Appendix I.
4 Process reward
The PRM takes in the problem and existing reasoning and outputs a numerical score evaluating its quality.
Here is an input-output example for the PRM:
Input:
(Problem) How many square units are in the region satisfying the inequalities
\$y $\ge$ |x|\$ and \$y $\le$ -|x|+3\$?
(Existing reasoning) Let’s start by analyzing the given inequalities. The next
step is to determine the points of intersection... Now we know the points of
intersection are...
Output:
(Reward) 0.765
5 Definition of MDP
Combining the above components, our MDP is defined as follows:
  • State: A low-dimensional vector generated from the LLM’s self-evaluation of the current reasoning step.
  • Action: One of five logic blocks that guide the next step reasoning.
  • Reward: A score from a PRM that evaluates the quality of the reasoning step after an action is taken.
6 Training process
Follow the MDP framework, we train the navigator with Deep Q-Learning algorithm. The training process is:
- For training episodes:
  - Randomly sample a Problem. Existing reasoning={}.
  - WHILE True:
    - Problem + Existing reasoning --[Self-evaluation]--> State vector
    - State vector --[MLP navigator]--> Logic block (action)
    - Problem + Existing reasoning --[Logic block]--> New reasoning
    - Existing reasoning <-- Existing reasoning + New reasoning
    - Problem + Existing reasoning --[PRM]--> Reward
    - Use the reward to train the MLP navigator
    - IF reach the answer: BREAK
7 Inference process
With the trained navigator, the inference process is:
- Problem. Existing reasoning={}. Answers={}.
- For self-consistency candidates number:
  - WHILE True:
    - Problem + Existing reasoning --[Self-evaluation]--> State vector
    - State vector --[MLP navigator]--> Logic block (action)
    - Problem + Existing reasoning --[Logic block]--> New reasoning
    - Existing reasoning <-- Existing reasoning + New reasoning
    - IF reach the answer: BREAK
  - Answers <-- Answers + New answer
- Answers --[Self-consistency]--> Final answer
8 Self-consistency evaluation
To enhance the robustness of our final answers, we employ a self-consistency mechanism. Since the final answers can be in different formats, we use sympy.parsing.latex library to parse the answers and determine if they are mathematically equivalent. We select the final answer using majority voting. If no majority exists, one answer is selected at random.
Here is an input-output example for self-consistency:
Input:
(Answers) {0.5} {0.7} {1/2}
Output:
(Final answer) 0.5
Appendix K Detailed Prompts
In this section, we provide the prompts for obtaining states and performing actions to ensure reproducibility. In the following prompt blocks, the text enclosed in braces "{}" denotes problem-specific content, such as intermediate reasoning steps, while the remaining texts serve as fixed templates.
K.1 Prompt for Obtaining States
We use the following prompt to extract a state vector from an intermediate reasoning step. This prompt guides the LLM in systematically evaluating the current step across multiple aspects. During our experiments, we observed that LLMs effectively identify these aspects and provide detailed scores for each aspect.
Listing 1: Prompts for obtaining states
1{Problem and reasoning steps}
2Please evaluate the current step from the following aspects.
3A) Correctness
4 A1: Correctness of modeling:
5 Whether the current step is correctly derived from the origin problem.
6 A2: Clarity for further reasoning:
7 Whether the current step is clearly presented, without ambiguity, to support further reasoning.
8 A3: Correctness of calculation:
9 Whether the numerical computation in the current step is performed correctly.
10B) Complexity
11 B1: Complexity to reach the final answer:
12 Whether it still requires complex reasoning or calculation to reach the final answer from the current step.
13 B2: Alternative methods in further reasoning:
14 Whether there exist multiple alternative methods to solve the problem in the current step.
15C) Completeness
16 C1: Closeness to the final solution:
17 Whether the current step is close enough to directly reach the final answer.
18 C2: Completeness within the step:
19 Whether all necessary elements within this specific step are known from the problem or previous steps.
20For each aspect, please score 1 for False, 2 for Unsure, and 3 for True, and score 0 if the current step does not involve this aspect. Please attach the reason for each score.
21Use the format A1 score=[SCORE] reason=[REASON]’.
22Only score the current reasoning step here, and DONOT conduct further reasoning.
K.2 Prompts for Actions
Reason one step: The prompt below is designed to conduct the action of “Reason one step”. We emphasize that it should focus on one step at a time, which better controls the output and prevents mistakes in long reasoning paths.
Listing 2: Prompts for action “Reason one step”
1Here is a problem and several reasoning steps.
2{Problem and previous steps}
3Please reason exactly ONE more step based on the current step here, and DONOT reason too many steps at once.
Decompose: A “Decompose” action consists of the following three prompts. First, we use prompt 3 to break down the current problem into multiple subtasks. Next, prompt 4 is applied sequentially to each subtask to complete its execution. Since the current subtask may depend on previous ones, the results of earlier subtasks are included (see line 4 of the prompt). Finally, the execution results are summarized using prompt 5, which captures the key steps and outcomes. Only the summarized result is utilized for subsequent reasoning actions.
Listing 3: Prompts for obtaining subtasks in action “Decompose”
1Here is a problem and several reasoning steps.
2{Problem and previous steps}
3Please decompose the current task into subtasks, where we can solve the original problem by combining these results of subtasks.
4Only provide subtasks decomposition here, and DONOT conduct specific reasoning or calculation.
5Use the format ’### Subtask1: subtask1’.
Listing 4: Prompts for executing subtasks in action “Decompose”
1Here is a problem and several reasoning steps.
2{Problem and reasoning steps before decomposition}
3For the next step, the task is decomposed into subtasks, here are the reasonings in the first few subtasks.
4{Executing results of previous subtasks}
5Please conduct the following Subtask{subtask_id} to continue the reasoning. text_decompose.
6DONOT conduct a more detailed decomposition for the subtask.
Listing 5: Prompts for summarize subtasks in action “Decompose”
1Here are a few detailed reasoning subtasks of a problem.
2{Executing results of subtasks}
3Please give a clear and concise summary of these subtasks, keeping the key reasoning and results in each subtask.
4Only provide the summary here, and DONOT conduct more reasoning or calculation.
Debate: A "Debate" action involves multiple rounds of question-answering. First, the LLM generates different plans for the task using prompt 6. Next, the plans are compared, and the most promising one is selected using prompt 7, mimicking how human experts debate and discuss to reach a solution. Based on the chosen plan, reasoning is advanced by one step through prompt 8. Similar to the “Decompose” action, only the results of the final one-step reasoning (output of prompt 8) are retained for subsequent reasoning processes.
Listing 6: Prompts for obtaining various plans in action “Debate”
1Here is a problem and several reasoning steps.
2{Problem and previous reasoning steps}
3Please propose three different alternative plans for solving the problem in the current step.
4Only provide plans here, and DONOT conduct specific reasoning or calculation.
5Use the format ’### Plan1: plan1’.
Listing 7: Prompts for analysing and comparing plans in action “Debate”
1Here is a problem and several reasoning steps.
2{Problem and previous reasoning steps}
3Currently, we have several alternative plans for solving the problem in the current step.
4{Generated Plans}
5Please review and compare these plans carefully, and tell which one is most promising for further reasoning. Only compare the plans here, and DONOT conduct further reasoning or calculation.
6Use the format The most promising plan is Plan[INDEX]: [REASON]’, where [INDEX] is an integer index of the plan and [REASON] is a detailed analysis.
Listing 8: Prompts for executing the plan in action “Debate”
1Here is a problem and several reasoning steps.
2{Problem and previous reasoning steps}
3For the next step, we have decided on the most promising plan:
4{Plan}
5Please reason **exactly one** more step according to the plan here, and DONOT reason too many steps at once.
Refine: We use prompt 9 to perform the “Refine” action, which instructs the LLM to review and improve the reasoning steps for clarity and correctness.
Listing 9: Prompts for action “Refine”
1Here is a problem and several reasoning steps
2{Problem and previous reasoning steps}
3Please check and refine the current thought here, and DONOT conduct further reasoning or calculation.
Terminate: The prompt for the “Terminate” action concludes the reasoning process by generating the final answer. The only variation lies in the output format, which is adapted to the specific requirements of each dataset.
Listing 10: Prompts for action “Terminate”
1Here is a problem and several reasoning steps
2{Problem and previous reasoning steps}
3
4## GSM8K
5Please generate the answer for the problem. Please end the answer with The answer is numerical_answer’.
6
7## MATH
[…]
9
10## MMLU-STEM and GPQA
11End the answer with The answer is (CHOICE)’.
12
13## StrategyQA
14Please generate the answer for the problem. At the end of your answer, conclude the answer with The answer is yes or The answer is no’.
Appendix L Use of LLMs
The authors used LLMs to aid or polish paper writing, but all content has been carefully reviewed by the author. The authors used LLMs for literature retrieval and discovery, but all related works have been carefully reviewed and organized by the author. The research ideation in this work was entirely completed by the author and does not involve the use of LLMs.

### v3

@@
 During the reasoning process, the trained RL navigator dynamically selects the suitable logic blocks and combines them into task-specific logical structures according to problem characteristics.
-Experiments across multiple reasoning benchmarks (AIME, MATH, GPQA, etc.) with multiple LLMs (GPT, Llama, Qwen, and DeepSeek) illustrate that RLoT outperforms established inference-time techniques by up to 13.4%.
+Experiments across multiple reasoning benchmarks (AIME, MATH, GPQA, etc.) with multiple LLMs (GPT, Llama, Qwen, and DeepSeek) illustrate that RLoT outperforms established inference-time techniques in most cases and improves up to 13.4% in challenging situations.
 Remarkably, with less than 3K parameters, our RL navigator is able to make sub-10B LLMs comparable to 100B-scale counterparts.
 Moreover, the RL navigator demonstrates strong transferability: a model trained on one specific LLM-task pair can effectively generalize to unseen LLMs and tasks.
-Our code is open-source at https://anonymous.4open.science/r/RL-LLM-Reasoning-1A30.
+Our code is open-source at https://github.com/tsinghua-fib-lab/RL-LLM-Reasoning.
 1 Introduction
@@
 Detailed A1: Correctness of modeling B1: Complexity to the final answer C1: Closeness to the final solution
-A2: Clarity for further reasoning
-B2: Alternative methods in
-further reasoning
-C2: Completeness within the step
+A2: Clarity for further reasoning B2: Alternative methods in further reasoning C2: Completeness within the step
 A3: Correctness of calculation
@@
 Within the MDP framework outlined above, given a specific LLM type and a kind of reasoning task, we train the navigator model.
+Under this formulation, the training process constitutes a standard RL problem within a discrete action space.
+Consequently, our framework is algorithm-agnostic, allowing for the employment of arbitrary off-the-shelf RL algorithms for training.
 To enhance the learning for challenging reasoning tasks, we extract hard questions from the training set of the target task, i.e., questions that the LLM cannot answer when directly prompted.
 Then, we use these problems for training the navigator model, from which we randomly select one in each episode and repeat it multiple times.
-We employ the Double-Dueling-DQN algorithm (Van Hasselt et al., 2016; Wang et al., 2016), and include the online and target network technique to improve the stability during training (Mnih et al., 2015).
 We illustrate the training and inference pipeline of RLoT in Figure 1.
@@
 These multi-round approaches facilitate comparison and voting across diverse reasoning paths, thereby enhancing the reasoning capability of LLMs in complex tasks.
-Implementation of RLoT.
+4.2 Training Details
 In our implementation, the navigator model is a simple three-layer multilayer perceptron (MLP) with the Dueling Network architecture (Wang et al., 2016).
 The model merely contains 2,566 parameters in total, where the lightweight design ensures efficient training and inference.
-Following existing works like CoT-SC, we perform multiple repeated inferences for each task and filter out the trails that do not meet self-consistency, enhancing the robustness of reasoning.
+We employ the Double-Dueling-DQN algorithm (Mnih et al., 2015; Van Hasselt et al., 2016; Wang et al., 2016) for optimization.
+By integrating Double Q-learning to mitigate value overestimation and the Dueling architecture to separate state and advantage representations, this algorithm significantly improves stability during the training process.
+We train the navigator model for 3,000 episodes with Qwen2.5-14B-Instruct on the MATH benchmark, where the learning curves presented in Figure 2 and Appendix A.1 indicate strong convergence.
+Furthermore, in Appendix A.2, we investigate the impact of employing alternative RL algorithms and extending training episodes, where the results consistently demonstrate similar convergence.
+Figure 2: Learning curves during RL training of navigator models.
 To obtain reward signals for actions, we utilize the Math-Shepherd as the PRM (Wang et al., 2024c).
@@
 ToT 13.33 50.00 73.80 91.35 36.60 82.62 70.45 59.74
-RLoT (ours) 23.33 60.00 76.7 92.87 44.64 85.06 79.04 65.95
+RLoT (ours) 23.33 60.00 76.70 92.87 44.64 85.06 79.04 65.95
 Llama3.1-8B-Instruct Direct QA 49.12 84.83 32.14 71.58 70.74 61.68
@@
 RLoT (ours) 20.00 70.00 77.36 93.86 54.02 88.23 82.68 69.45
-DeepSeek-R1-
-Distill-Qwen-7B
-Direct QA 46.67 60.00 92.27 95.74 54.47 83.61 79.48 73.18
+DeepSeek-R1- Distill-Qwen-7B Direct QA 46.67 60.00 92.27 95.74 54.47 83.61 79.48 73.18
 Zero-shot CoT 53.33 62.50 91.48 95.20 53.13 85.28 78.89 74.26
@@
 4.3 Overall Performance
-We train the navigator model with Qwen2.5-14B-Instruct on the MATH benchmark, where the learning curve shown in Appendix A indicates good convergence.
-We use the obtained navigator model to enhance the reasoning of multiple LLMs across different reasoning tasks and present the results in Table 2.
+We use the obtained navigator model to enhance the reasoning of multiple LLMs across different reasoning tasks.
+Following existing works like CoT-SC, we perform multiple repeated inferences for each task and filter out the trails that do not meet self-consistency, enhancing the robustness of reasoning, where the results are presented in Table 2.
 For the baseline performance, we prioritize using the results reported in the official technical reports of each LLM (Yang et al., 2024a; Dubey et al., 2024; Hurst et al., 2024), if available.
@@
 Task MATH GPQA StrategyQA
-2-step pattern Reason-Refine Reason-Refine Reason-Debate
+Instance number 5000 448 687
+Major 2-step pattern Reason-Refine Reason-Refine Reason-Debate
 Reason-Decompose Reason-Debate Reason-Refine
-3-step pattern Decompose-Refine-Reason Reason-Refine-Debate Reason-Decompose-Debate
+Major 3-step pattern Decompose-Refine-Reason Reason-Refine-Debate Reason-Decompose-Debate
 Reason-Refine-Debate Reason-Decompose-Refine Reason-Refine-Debate
@@
 These typical reasoning patterns exhibit strong interpretability, further validating the capability of RLoT to generate task-specific logical structures that enhance LLM reasoning.
-For better understanding, we provide an example in Appendix B to illustrate how these logical structures empower the correct solution to a problem.
+For better understanding, we provide an example in Appendix B.1 to illustrate how these logical structures empower the correct solution to a problem.
 4.7 Ablation Studies, Analyses, and Discussions
-To better justify our designs, we conducted a series of ablation studies, analyses, and discussions.
+To better justify our designs, we verify the effectiveness of our logic block designs with ablation studies.
+To examine the impact of each block, we remove each of them and train a navigator, where results in Table 6 confirm that all blocks are effective.
+Table 6: Ablation study results for each logic block.
+LLM Ablation MATH GPQA StrategyQA Average
+Qwen2.5-Instruct-7B w/o Decompose 75.42 31.92 77.00 61.45
+w/o Debate 74.02 36.61 77.58 62.74
+w/o Refine 75.76 41.29 72.93 63.33
+Full RLoT 76.70 44.64 79.04 66.79
+GPT-4o-mini w/o Decompose 76.08 41.74 79.33 65.72
+w/o Debate 74.68 45.31 78.75 66.25
+w/o Refine 76.44 51.56 74.38 67.46
+Full RLoT 77.36 54.02 82.68 71.35
+Also, we conducted a series of analyses, and discussions.
+We quantify the contribution of each logic block in Appendix I.
 We discuss the reliability of the self-evaluation state design in Appendix F.
@@
 We illustrate the role of the process reward model (PRM) in training the navigator in Appendix H.
-Also, we verify the effectiveness of our logic block designs and quantify the contribution of each logic block in Appendix I.
 5 Related Works
@@
 For Reproducibility, we describe the general experimental settings in Section 4; we list the implementation details in Appendix J; and our source code are anonymously open source at https://anonymous.4open.science/r/RL-LLM-Reasoning-1A30.
-Appendix A Learning Curves
+Appendix A Training Details
+A.1 Learning Curves
 In Figure 3, we show the learning curves during RL training of all navigator models used in experiments in the main text.
 We use a sliding window averaging to smooth the reward, and from the learning curves, we observe good convergence of the RL training.
-Figure 2: Learning curves during RL training of all navigator models.
+Figure 3: Learning curves during RL training of navigator models.
+A.2 Analysis
+In Figure 4, we extend the training duration to 5,000 episodes.
+We observe that the model essentially reaches convergence around 3,000 episodes, with only marginal performance gains achieved thereafter.
+Additionally, we explore training the navigator model using the Double-DQN algorithm.
+While the training stability decreases slightly due to the absence of the Dueling network architecture, the model ultimately converges to a performance level comparable to that of the Double-Dueling-DQN.
+Figure 4: 5k episodes learning curves during RL training with different algorithms.
 Appendix B Case Study
+B.1 Correct Cases
 Figure 5: A case study comparing few-shot CoT and RLoT on a representative problem in the MATH dataset. Results are both generated by Qwen2.5-14B-Instruct.
@@
 By avoiding unnecessary geometric division, RLoT achieves the correct answer.
+B.2 Failure Cases
+To provide a comprehensive understanding of the boundaries of RLoT, we also analyze typical failure cases observed during our experiments.
+Common failures can be categorized into two primary modes: (1) Navigator input errors caused by self-evaluation noise, and (2) Navigator output errors caused by suboptimal action selection.
+As shown in Figure 6, failure happens due to self-evaluation noise.
+The robustness of the navigator depends partly on the accuracy of the state representation derived from the LLM self-evaluation.
+In this case, the LLM generates an incorrect equation (wrongly deriving 3q=24p363q=24p-36) but fails to identify this error during the self-evaluation phase, hallucinating a "True" status for correctness of modeling.
+This noise in the state input misleads the navigator: instead of selecting the necessary "Refine" action to correct the mistake, the Navigator perceives the current step as valid and selects "Debate" to explore alternative plans.
+Figure 6: A failure case study that is caused by wrong state self-evaluation.
+As shown in Figure 7, failure happens due to action selection error, where the navigator selects a suboptimal action despite receiving accurate state information.
+In this case, the LLM derives a correct intermediate result ($\cos(2\pi/7)+\dots=-1/2$) and correctly evaluates it as "True".
+However, the navigator triggers a "Refine" action instead of proceeding with reasoning or termination.
+This unnecessary intervention forces the LLM to modify a correct step, leading to "over-correction" where valid reasoning is replaced with an erroneous identity, ultimately causing the solution to fail.
+Figure 7: A failure case study that is caused by wrong action selection.
+While these failure modes exist, our extensive quantitative results demonstrate that RLoT achieves performance gains over baselines, indicating that these failure cases represent a minority of the inference trajectories.
+In our design, the RL training process effectively optimizes the navigator to minimize action selection errors, and the structured self-evaluation prompt is designed to mitigate state noise errors as much as possible.
 Appendix C Computational Overhead and Latency
@@
 RLoT (ours) 26529 4778 19513 3973 3483 1575 3069 1048 4372 1550 1882 710 1612 673 8637 2044
-DeepSeek-R1-
-Distill-Qwen-7B
-Direct QA 879 11793 597 4686 541 4960 533 1364 808 5462 521 2094 436 1268 616 4518
+DeepSeek-R1- Distill-Qwen-7B Direct QA 879 11793 597 4686 541 4960 533 1364 808 5462 521 2094 436 1268 616 4518
 Zero-shot CoT 906 12129 640 5140 565 4697 547 1381 822 5443 574 2012 450 1344 643 4592
@@
 Fixed reasoning patterns. Baselines from this category adopts a fixed workflow to obtain the answer, which
+  • Self-Refine (Madaan et al., 2023): This method adds a simple refine step after every reasoning step to correct potential mistakes.
+  • Least-to-Most (Zhou et al., 2023): Ask LLMs to break down a complex problem into a series of simpler subproblems and then solve them in sequence.
+  • SelfCheck (Miao et al., 2024): Utilize LLMs to check their own outputs and use the results of these checks to improve question-answering performance by conducting weighted voting on multiple solutions to the question.
+Here we adopt voting with 4 candidates.
   • DeAR (Xue et al., 2024): The problem is consequently decomposed, analyzed and refined in this workflow. The workflow follows a recurrent manner to a detailed decomposition of original problem.
-  • Self-Refine (Madaan et al., 2023): This method adds a simple refine step after every reasoning step to correct potential mistakes.
 Tree search.Tree-based models divide the reasoning task into sub-steps, and searches for a best path to the final answer.
@@
 These patterns are continuously updated during inference time.
-Table 9: Extended baseline comparisons. The bold numbers indicate the best performance in each group of experiments, and the underlined numbers indicate the best baseline method.
-LLM Method GSM8K GPQA StrategyQA Average
-Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
-DeAR (Xue et al., 2024) 88.38 36.16 71.13 65.22
-Self-Refine (Madaan et al., 2023) 88.55 32.59 73.94 65.03
-Tree search One-step-greedy 89.08 32.59 73.07 64.91
-Basic MCTS 90.07 34.82 75.69 66.86
-Litesearch (Wang et al., 2024a) 89.76 33.04 75.84 66.21
-Q* (Wang et al., 2024b) 91.51 34.60 77.00 67.70
-rStar (Qi et al., 2024) 91.96 38.62 77.15 69.24
-AFlow (Zhang et al., 2024b) 91.36 37.72 76.56 68.55
-DSPy) (Khattab et al., 2023) 88.32 34.15 74.09 65.52
-Graph of Thoughts (GoT) (Besta et al., 2024) 88.96 33.04 75.42 65.81
-Buffer of Thoughts (BoT) (Yang et al., 2024d) 92.35 34.51 74.33 67.06
-RLoT (ours) 92.87 44.64 79.04 72.18
+Table 10: Extended baseline comparisons.
+The bold numbers indicate the best performance in each group of experiments, and the underlined numbers indicate the best baseline method.
+All results are averaged across 5 repeated runs with standard deviation.
+LLM Method MATH-500 GSM8K GPQA StrategyQA Average
+Qwen2.5-Instruct-7B DirectQA 74.96±1.37 90.95±1.06 31.16±0.23 68.53±0.53 66.40±0.46
+Self-Refine 72.32±1.20 88.39±0.38 32.23±0.61 74.06±0.79 66.75±0.33
+Least-to-Most 73.28±0.84 89.10±0.82 31.79±0.54 73.60±0.64 66.94±0.25
+SelfCheck 74.36±1.60 90.24±1.09 33.84±0.61 72.87±0.77 67.83±0.35
+DeAR 69.92±0.88 87.84±0.54 35.71±0.37 71.03±0.85 66.13±0.39
+Tree search One-step-greedy 71.24±1.43 89.49±0.99 32.63±0.50 72.43±0.56 66.45±0.54
+Basic MCTS 72.40±0.79 89.51±0.72 35.13±0.59 75.72±0.80 68.19±0.19
+Litesearch 72.52±1.29 89.86±0.86 33.17±0.30 76.04±0.65 67.90±0.52
+Q* 73.76±0.81 92.28±0.61 34.38±0.76 76.80±0.99 69.30±0.50
+rStar 75.88±1.02 92.05±0.89 38.17±0.32 77.09±0.83 70.80±0.68
+AFlow 75.72±1.16 90.95±0.98 37.37±0.63 77.09±0.56 70.28±0.13
+DSPy 72.48±0.86 88.75±0.77 34.11±0.68 73.89±0.94 67.31±0.49
+Graph of Thoughts (GoT) 73.24±1.06 88.76±1.07 33.26±0.66 75.69±0.91 67.74±0.39
+Buffer of Thoughts (BoT) 75.16±0.74 91.98±0.64 34.91±0.58 74.38±0.44 69.11±0.32
+RLoT (ours) 76.88±1.35 92.95±0.63 44.78±0.71 79.56±0.64 73.54±0.18
 As shown in Table 10,we compare these methods on representative datasets, GSM8K, GPQA, and StrategyQA, which covers three different domains.
@@
 By flexibly organizing logic blocks, our design can outperform these baselines. To compare the efficiency and effectiveness, we show the token consumption per question for representative baselines:
-Table 10: Token consumption and performance scores per question for representative baselines.
+Table 11: Token consumption, solving time (minutes) per question, and performance scores for representative baselines.
 LLM Method GSM8K GPQA StrategyQA Average
-Score Token Score Token Score Token Score Token
-Qwen2.5-Instruct-7B DirectQA 91.58 240 31.25 705 68.85 301 63.89 415
-Self-Refine 88.55 678 32.59 1784 73.94 939 65.03 1134
-rStar 91.96 2636 38.62 4127 77.15 2987 69.24 3483
-RLoT (ours) 92.87 1759 44.64 3001 79.04 2341 72.18 2367
+Score Token Time Score Token Time Score Token Time Score Token Time
+Qwen2.5-Instruct-7B DirectQA 91.58 240 0.05 31.25 705 0.17 68.85 301 0.05 63.89 415 0.09
+Self-Refine 88.55 678 0.12 32.59 1784 0.41 73.94 939 0.16 65.03 1134 0.23
+Litesearch 89.76 912 0.20 33.04 2139 0.55 75.84 1228 0.24 66.21 1426 0.33
+rStar 91.96 2636 0.77 38.62 4127 1.44 77.15 2987 0.82 69.24 3483 1.01
+RLoT (ours) 92.87 1759 0.61 44.64 3001 0.95 79.04 2341 0.53 72.18 2367 0.70
 Unlike search-based methods that require multiple explorations, our design directly generates task-specific logical trajectories, allowing for better performance with a similar or lower cost.
@@
 Meanwhile, with substantially lower computational consumption, it reaches comparable performance to some fine-tuning methods that require modification on the parameters of LLMs.
-Appendix F Analyses on the Reliability of Self-evaluation State
+Appendix F Analyses on the Self-evaluation State
+F.1 Reliability
 Previous work shows that LLMs can evaluate and correct their own outputs (Han et al., 2024).
@@
 Modeling Calculation
-Mark has a garden with
-flowers. He planted plants
-of three different colors in it.
-Ten of them are yellow,
-and there are 80% more of
-those in purple. There are only
-25% as many green flowers as
-there are yellow and purple
-flowers. How many flowers
-does Mark have in his garden?
+Mark has a garden with flowers.
+He planted plants of three different colors in it.
+Ten of them are yellow, and there are 80% more of those in purple.
+There are only 25% as many green flowers as there are yellow and purple flowers.
+How many flowers does Mark have in his garden?
 In this step, we aim at calculate the number
@@
 Based on the validation, we consider our structured self-evaluation provides a sufficiently reliable state representation for training the navigator.
+F.2 Noise Impact
+To investigate the navigator’s reliance on accurate state perceptions, we introduced synthetic noise into the self-evaluation state vectors during inference.
+As shown in Table 14, replacing the state with random values results in a significant performance drop, falling even below DirectQA.
+This confirms that the navigator actively utilizes the state information for decision-making rather than following a fixed, state-agnostic policy.
+However, the framework exhibits commendable robustness.
+Under extreme conditions with 50% noise (where half of the state bits are flipped), RLoT still maintains an average accuracy of 64.99%, being comparable to the DirectQA baseline.
+As the noise decreases to 30%, performance recovers rapidly to 69.11%.
+Table 14: Impact of noise in self-evaluation states.
+LLM State GSM8K GPQA StrategyQA Average
+Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
+Random 89.76 28.13 65.50 61.13
+50% noise 91.74 34.82 68.41 64.99
+30% noise 92.27 43.75 71.32 69.11
+Normal 92.87 44.64 79.04 72.18
+Meanwhile, to minimize potential noise inherent in self-evaluation, we deliberately designed the evaluation criteria as a binary classification task rather than a multi-grade continuous scoring.
+Continuous or fine-grained scoring (e.g., 1-10 scales) requires precise calibration, which is challenging for LLMs.
+In contrast, a binary standard simplifies the decision boundary, significantly reducing the cognitive load on the LLM.
+This design choice makes the evaluation process more robust, ensuring that the state vector remains a stable and reliable signal for the navigator even with less capable LLMs.
+F.3 Alternative Designs
 Furthermore, we explore alternative designs of self-evaluation:
@@
 LLM Method GSM8K GPQA StrategyQA Average
-Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 66.62
-Repeated refine 87.72 32.37 72.34 66.65
-Repeated debate 89.99 33.04 72.05 67.51
-Fixed logic sequence 88.38 36.16 71.13 67.47
-Supervise-trained navigator 89.84 36.83 70.31 67.94
-LLM as navigator 89.16 37.95 74.38 69.17
-RL-trained navigator (ours) 92.87 44.64 79.04 73.34
+Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
+Repeated refine 87.72 32.37 72.34 64.14
+Repeated debate 89.99 33.04 72.05 65.03
+Fixed logic sequence 88.38 36.16 71.13 65.22
+Supervise-trained navigator 89.84 36.83 70.31 65.66
+LLM as navigator 89.16 37.95 74.38 67.16
+RL-trained navigator (ours) 92.87 44.64 79.04 72.18
 The results in Table 16 show that these methods are worse than RLoT:
@@
 These show the importance of using RL for our navigator to flexibly select logic blocks.
-Appendix H Analyses on the Role of Process Reward Model (PRM)
-As literature (Zhang et al., 2025) indicates that math-trained PRMs perform well on MMLU-STEM, and the performance of PRMs in math and STEM is generally correlated.
-Our navigator, which is trained with a math PRM, achieves strong performance on GPQA (STEM).
-This also reflects the generalization capability of PRM across domains.
-We also test an ORM to train the navigator.
+Appendix H Analyses on the Process Reward Model (PRM)
+H.1 The Role of PRM
+To investigation the role of PRM, we also test an ORM to train the navigator.
 Table 17: Comparison of RLoT using an Outcome Reward Model (ORM) vs. a Process Reward Model (PRM).
 LLM Method GSM8K GPQA StrategyQA Average
-Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 66.62
-RLoT with ORM 91.96 41.52 73.94 70.71
-RLoT with PRM (ours) 92.87 44.64 79.04 73.34
+Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
+RLoT with ORM 91.96 41.52 73.94 69.14
+RLoT with PRM (ours) 92.87 44.64 79.04 72.18
 The results in Table 17 show that the PRM outperforms the ORM, indicating that the step reward signal from PRM is crucial.
+H.2 Generalization Capability of PRM
+Recent studies suggest that process reward models (PRMs) trained on mathematical data exhibit strong transferability to related domains such as STEM (Zhang et al., 2025).
+To empirically verify this and justify our reward design, we conducted a PRM calibration analysis.
+We calculated the correlation between the intermediate PRM scores assigned at various reasoning steps and the correctness of the final answer.
+Table 18: PRM calibration.
+LLM Correlation with Correctness MATH GPQA StrategyQA
+Qwen2.5-Instruct-7B Step1 PRM 0.199 0.184 0.157
+Step2 PRM 0.222 0.213 0.195
+Step3 PRM 0.236 0.215 0.202
+Step4 PRM 0.240 0.228 0.218
+Average 0.224 0.210 0.193
+As presented in Table 18, the results reveal two insights.
+First, there is a consistent positive correlation between intermediate PRM scores and the final outcome across all tested benchmarks.
+This confirms that the PRM effectively gauges the quality of intermediate steps, validating its suitability as a dense reward signal for training the navigator.
+Second, although the PRM was trained primarily on mathematical data, its scores exhibit a notable correlation with solution correctness on out-of-domain tasks, including GPQA (STEM) and StrategyQA (commonsense).
+This demonstrates the PRM’s intrinsic ability to generalize its verification logic across different domains.
+This calibration analysis explains the underlying mechanism of our method’s transferability.
+Because the math-trained PRM provides reliable feedback even on unseen tasks like GPQA and StrategyQA, the navigator trained with these rewards can successfully learn effective reasoning policies across diverse domains.
+H.3 Impact of PRM Quality
 Further, to explore the impact of PRM quality, we test:
@@
 Appendix I Analyses on the Design of Logic Blocks
+I.1 Effect of Logic Block Number
 We first check the test-time scaling for the logic blocks. Results in Figure 8 show that a longer sequence of logic blocks brings a performance gain.
 Figure 8: Test-time scaling with different number of logic blocks.
-To examine the impact of each block, we remove each of them and train a navigator:
-Table 17: Ablation study results for each logic block.
-LLM Ablation MATH GPQA StrategyQA Average
-Qwen2.5-Instruct-7B w/o Decompose 75.42 31.92 77.00 61.45
-w/o Debate 74.02 36.61 77.58 62.74
-w/o Refine 75.76 41.29 72.93 63.33
-Full RLoT 76.70 44.64 79.04 66.79
-GPT-4o-mini w/o Decompose 76.08 41.74 79.33 65.72
-w/o Debate 74.68 45.31 78.75 66.25
-w/o Refine 76.44 51.56 74.38 67.46
-Full RLoT 77.36 54.02 82.68 71.35
-The ablation results in Table 17 confirm that all blocks are effective.
+I.2 Role of Different Logic Blocks on Different Datasets
 We also analyze the impact of different logic blocks on different datasets.
@@
 Figure 9: Contribution of each logic block in the reasoning sequence.
+I.3 Alternative Blocks
 Further, we test some possible new blocks:
@@
 "Decompose," "Debate," and "Refine" are the core actions, supplemented by "Reason one step" and "Terminate" to ensure complete reasoning flows.
-For instance, "Decompose" reflects the widely studied divide-and-conquer approach, while "Debate" aligns with strategies involving comparative evaluation (Wang & Chiew, 2010).
+For instance, "Decompose" reflects the widely studied divide-and-conquer approach, while "Debate" aligns with strategies involving comparative evaluation (Wang and Chiew, 2010).
 Also, they are empirically effective.
@@
 This allows flexible composition of logic blocks.
-Third, here is an input-output example for the ’Debate’ block (as Appendix B):
+Third, here is an input-output example for the ’Debate’ block (as Appendix B.1):
 Input
@@
 (Final answer) 0.5
-Appendix K Detailed Prompts
+Appendix K Prompts
 In this section, we provide the prompts for obtaining states and performing actions to ensure reproducibility.
@@
 4{Executing results of previous subtasks}
-5Please conduct the following Subtask{subtask_id} to continue the reasoning. text_decompose.
+5Please conduct the following Subtask{subtask_id} to continue the reasoning.
 6DONOT conduct a more detailed decomposition for the subtask.
@@
 14Please generate the answer for the problem. At the end of your answer, conclude the answer with The answer is yes or The answer is no’.
+K.3 Sensitive to action-prompts phrasing.
+To evaluate whether the performance of RLoT is sensitive to the specific wording of the action prompts, we conducted a robustness test by rephrasing the natural language instructions for the Decompose, Debate, and Refine actions.
+Decompose: We rephrase the prompts for the action of Decompose as:
+Listing 11: Prompts for obtaining subtasks in action “Decompose”
+1This is the problem and the reasoning progress so far.
+2{Problem and previous steps}
+3Break down the current problem into smaller, manageable sub-problems.
+4Solving these sub-problems sequentially should lead to the solution of the original question.
+5Only provide subtasks decomposition here, and DONOT conduct specific reasoning or calculation.
+6Format: ’### Subtask1: subtask1’.
+Listing 12: Prompts for executing subtasks in action “Decompose”
+1This is the problem context and the steps taken.
+2{Problem and reasoning steps before decomposition}
+3We have divided the main task into several parts. Below is the reasoning progress for the initial parts.
+4{Executing results of previous subtasks}
+5Now, please execute the specific instruction for Subtask{subtask_id} to advance the solution.
+6DONOT conduct a more detailed decomposition for the subtask.
+Listing 13: Prompts for summarize subtasks in action “Decompose”
+1Below are the detailed execution steps for the problems subtasks.
+2{Executing results of subtasks}
+3Synthesize the steps above into a brief and coherent overview, ensuring that the essential logic and final outcomes of each subtask are preserved.
+4Only provide the summary here, and DONOT conduct more reasoning or calculation.
+Debate: We rephrase the prompts for the action of Debate as:
+Listing 14: Prompts for obtaining various plans in action “Debate”
+1This is a problem and existing reasoning steps.
+2{Problem and previous reasoning steps}
+3Formulate three distinct strategies to address the problem from the current state.
+4Only provide plans here, and DONOT conduct specific reasoning or calculation.
+5Use the format ’### Plan1: plan1’.
+Listing 15: Prompts for analysing and comparing plans in action “Debate”
+1This is the problem context and the steps taken.
+2{Problem and previous reasoning steps}
+3We have generated a set of potential strategies to address the problem at this stage. {Generated Plans}
+4Critically evaluate these options and identify the most viable strategy to advance the reasoning.
+5Only compare the plans here, and DONOT conduct further reasoning or calculation.
+6Use the format The most promising plan is Plan[INDEX]: [REASON]’, where [INDEX] is an integer index of the plan and [REASON] is a detailed analysis.
+Listing 16: Prompts for executing the plan in action “Debate”
+1Here is a problem and several reasoning steps.
+2{Problem and previous reasoning steps}
+3We have selected the optimal strategy to proceed: {Plan}
+4Carry out the immediate next logical step following this strategy, and DONOT reason too many steps at once.
+Refine: We rephrase the prompts for the action of Refine as:
+Listing 17: Prompts for action “Refine”
+1Here is a problem and several reasoning steps
+2{Problem and previous reasoning steps}
+3Please examine the current reasoning for potential improvements and revise it accordingly, and DONOT conduct further reasoning or calculation.
+Table 23: Impact of prompt rephrasing.
+LLM Method GSM8K GPQA StrategyQA Average
+Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
+RLoT (original) 92.87 44.64 79.04 72.18
+RLoT (rephrased) 92.49 45.09 79.91 72.42
+The results, presented in Table 23, demonstrate that RLoT exhibits high robustness to prompt variations.
+The model with rephrased prompts achieves an average accuracy very similar to the original one.
+This consistency indicates that the effectiveness of RLoT stems from the adaptive logical structures constructed by the navigator, rather than overfitting to specific phrasing or reliance on extensive prompt engineering.
 Appendix L Use of LLMs